OpenStudy (anonymous):

Estimate the following product using a power of ten. 9.1 × 43.27

OpenStudy (tkhunny):

So, 9.1 isn't too far from 10. And, 43.24 is around 4*10 or 5*10, leaning toward 4*10 10 *(4*10) = 10*10*4 = 100*4 = 400 or so.

OpenStudy (anonymous):

3.93757*10^5

OpenStudy (anonymous):

ok @jenni4r

OpenStudy (tkhunny):

Except that's an exact value with an incorrect exponent on the 10. It's also no attempt at all at estimating.

OpenStudy (anonymous):

@hartnn who is right?

OpenStudy (anonymous):

@Aperogalics who is right???

OpenStudy (tkhunny):

What does the question want? If an estimate is wanted, I showed one way to do that. There are other ways. If an exact answer is required, then 9.1 * 43.27 = 393.757 = 3.93757 * 10^2

OpenStudy (anonymous):

@Hares333 @tkhunny is correct but @tkhunny this works for liitle values only if we take larger value it will varies a lot in product:)

OpenStudy (tkhunny):

Absolutely there is significant variation. It is a very vague problem statement and deserves a response with a lot of leeway. Personally, I'd prefer a better problem statement. :-)

OpenStudy (anonymous):

It is a selection problem. The choices for the answer are 360, 432.7, 4,327, and 3,600.

OpenStudy (tkhunny):

And thus we see the value of getting the WHOLE problem statement. My original guess was magically pretty close to what was wanted. 1) My very rough estimate was sufficient to eliminate 4327 and 3600. Wrong orders of magnitude. 2) Since my estimate was HIGH (I went up to 10) and LOW (I used 40 rather than 50), I'm struggling with eliminating either 360 or 432.7 with only this evidence. 3) Deliberately low 9*40 = 360 ---- 432.7 is starting to look pretty good. 4) Deliberately high 10*50 = 500 - meh. 5) It's still not perfectly clear what is wanted.